OpenStudy (anonymous):

You roll two dice. What is the probability that the sum of the dice is even or the sum of the dice is less than 6? Make sure you only account for dice sums that are both even and less than 6 once. Write your answer in the form of a fraction, using the slash ( / ) key. Reduce fractions to their lowest terms.

OpenStudy (sriram):

less than 6 means 6 is not allowed right?

OpenStudy (sriram):

gettin less than 6 by 2[4+3+2+1]=20ways

OpenStudy (amistre64):

make a 6x6 table

OpenStudy (sriram):

common ways=4

OpenStudy (sriram):

hence probability=[18+20-4]/36=34/36

OpenStudy (amistre64):

sum is even sum is less than 6 1 2 3 4 5 6 1 x x x x 2 x x x 3 x x 4 x 5 6 looks to be 10 out of 36; or 10/36, simplify as needed

OpenStudy (sriram):

but even 6,6 will be counted

OpenStudy (amistre64):

hmmm..... yeah, evens above six count as well dont they :)

OpenStudy (amistre64):

in other words, count all the evens; and count all the less then 6s ...

OpenStudy (amistre64):

1 2 3 4 5 6 1 x x x x x 2 x x x x x 3 x x x x 4 x x x x 5 x x x 6 x x x 24 out of 10 perhaps :)

OpenStudy (amistre64):

and by 10 i mean 36 lol